OpenStudy (anonymous):

48÷2(9+3) = ? Is it 2 or 288? Justify your answer please, it's bugging me...

OpenStudy (anonymous):

ok first you need to add the numbers inside the parentheses....that's the rule. Then using parentheses you multiply 2 by 12, which is 24. 48 divided by 24 is 2.

OpenStudy (anonymous):

i think that it is 2 because you go from left to right

OpenStudy (anonymous):

My question is basically, is the question viewed as: (48/2)*(9+3) or 48/(2*(9+3))?

OpenStudy (anonymous):

it's 48 / (2*(9+3))

OpenStudy (anonymous):

Sorry Gang, it's (48/2) x (9+3) as it is written.

OpenStudy (anonymous):

\[48\div 2(9+3) = 48 \div 2 * (9+3) = 48\div 2 * 12 = \frac{48}{2}*12\] \[=24*12 = 36\]

OpenStudy (anonymous):

"24 * 12 = 36" *Awwwwwwwwwkward*

OpenStudy (anonymous):

Err yes, 288